If the puck has a mass of 0.1kg and Dan can only apply 200 Newton’s of force, how fast with the puck accelerate
1
Expert's answer
2020-03-30T07:53:23-0400
"a=\frac{F}{m}=\frac{200}{0.1}=2000\frac{m}{s^2}"